# Can LLMs Understand Social Norms in Autonomous Driving Games?

### v1

@@ Abstract
Social norm is defined as a shared standard of acceptable behavior in a society. The emergence of social norms fosters coordination among agents without any hard-coded rules, which is crucial for the large-scale deployment of AVs in an intelligent transportation system. This paper explores the application of LLMs in understanding and modeling social norms in autonomous driving games. We introduce LLMs into autonomous driving games as intelligent agents who make decisions according to text prompts. These agents are referred to as LLM-based agents. Our framework involves LLM-based agents playing Markov games in a multi-agent system (MAS), allowing us to investigate the emergence of social norms among individual agents. We aim to identify social norms by designing prompts and utilizing LLMs on textual information related to the environment setup and the observations of LLM-based agents. Using the OpenAI Chat API powered by GPT-4.0, we conduct experiments to simulate interactions and evaluate the performance of LLM-based agents in two driving scenarios: unsignalized intersection and highway platoon. The results show that LLM-based agents can handle dynamically changing environments in Markov games, and social norms evolve among LLM-based agents in both scenarios. In the intersection game, LLM-based agents tend to adopt a conservative driving policy when facing a potential car crash. The advantage of LLM-based agents in games lies in their strong operability and analyzability, which facilitate experimental design.
@@ I-B Contributions of this paper
This work focuses on how LLMs can guide agents in autonomous driving games. Our study utilizes the OpenAI Chat API, powered by GPT-4.0, to conduct experiments in simulated driving environments. The contributions include:
@@
    We propose a framework where LLM-guided agents play Markov games in a multi-agent system (MAS) in order to investigate whether social norms emerge among individual agents.
@@
    We design prompts to simulate interactions in an autonomous driving game. The prompts consist of “system” messages outlining the general set-up (e.g., the road layout) and “user” messages regarding the observations and decision making of LLM-guided agents.
@@
    We apply the proposed framework to two traffic scenarios: the unsignalized intersection and highway platoon, to examine social norms formed by autonomoud vehicles (AVs).
@@ II-A Markov Game
Markov games, also known as stochastic games, are a mathematical framework that generalizes both Markov decision processes (MDPs) and game theory to model dynamic interactions among rational players in uncertain environments. In its general form, it is defined by a states set $\mathcal{S}$ and action sets $\mathcal{A}_{1},...,\mathcal{A}_{n}$ for each corresponding agent in the environment. Each agent also has a corresponding reward function $\mathcal{R}i:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ for agent $i$. The goal of the Markov Game is to find an optimal policy for each agent.
In this paper, we adopt the general form of Markov Game with a multi-agent system. We want to observe the behavior and decision-making of LLMs when playing Markov Game as agents. We use $\mathcal{S}$ to denote the environment set-up which is fed to system agent, $\mathcal{A}_{i}$ to denote the action set of agent $i$, $\mathcal{O}_{i}$ to denote the observation set of agent $i$, $\mathcal{R}_{i}$ to denote the current reward of agent $i$. The actions of $\mathcal{A}_{i}$ are made by LLM(agent $i$) with the input of $\mathcal{O}_{i}$ and part of $\mathcal{S}$. The reward $\mathcal{R}_{i}$ of agent $i$ is output by LLM.
The optimal policy of agent $i$ is decided solely by LLMs, so we can investigate whether social norms are formed within the context of the above Markov Game.
II-B Social Norms
Social norms represent the collective expectations regarding appropriate conduct within groups. They encompass both the unwritten agreements that dictate how society members should act, and the formalized regulations and laws that are established. Social norms in car driving are the unwritten rules that guide driver behavior on the road, focusing on safety, courtesy, and communication. These include obeying traffic laws, signaling turns, maintaining safe distances, yielding when necessary, using headlights properly, acknowledging other drivers, and staying in the correct lane.
In this paper, our primary focus is on examining whether the agents could develop social norms, including avoiding crashes and yielding, in two specific scenarios.
II-C Large Language Models
@@
A large language model (LLM) is a type of artificial intelligence algorithm designed to understand, generate, and interact with human language at a large scale. These models are trained on vast amounts of text data, learning patterns, vocabulary, grammar, and even nuances of language such as context, tone, and implications. They use a specific architecture known as Transformer, which allows them to efficiently process sequences of words and predict the next word in a sentence, understand the meaning of a text, or generate new text that follows a given prompt.
LLMs are highly versatile AI tools with applications spanning multiple industries and fields. They excel in tasks such as content creation, customer support, education, language translation, software development, legal documentation, healthcare research, optimization, sentiment analysis, and accessibility technologies. In this paper, we examine the application of LLM in Markov games and autonomous vehicles to see if they will make appropriate decisions and form the social norms.
II-D Scenario 1: Unsignalized Intersection
The environment set $\mathcal{S}$ of a unsignalized intersection is characterized by two intersecting thoroughfares, with Road 1 extending from the West to the East and Road 2 extending from the North to the South. Green and red cars symbolize the agents traveling on Road 1 and Road 2, respectively. Agents aim to cross the intersection with action set $\mathcal{A}_{i}$ for agent $i$. Meanwhile, there will be some background vehicles that have predetermined driving strategies and do not develop any driving policy in Markov games. The background vehicles are included in $\mathcal{S}$. Since there is no traffic signals, agents need to observe the position of other vehicles and make appropriate decisions to avoid accidents. The observation is included in $\mathcal{O}_{i}$ for agent $i$.
Table I shows the set-up of agent $i$ at the beginning of time $t$. In our set-up, agent $i$’s location is which cell it is current at. The observation $o_{i}$ is a subset of current state space $\mathcal{S}$. Each agent can observe its own state $s_{i}$ in its entirety - both the location and the cumulative reward, but could only see the locations (not the cumulative rewards) of all other cars. The effect(i.e. the reward and state update) caused by the action $a_{i}[t]$ will not manifest until the beginning of the next time step $t+1$. In other words, the reward $r_{i}[t]$ reflects on the cumulative reward of $t+1$. The action $a_{i}[t]$ reflects on the location at $t+1$.
$s$ [${}_{i}[t]\in$ $\mathcal{S}$ Location and cumulative reward
$o$ [${}_{i}[t]\in$ $\mathcal{O}$ Observation of the current environment
$a$ [${}_{i}[t]\in$ $\mathcal{A}$ Go: move forward by one cell; Stop: no move
[…]
TABLE I: Intersection set-up
Figure 1: Unsignalized Intersection
II-E Scenario 2: Highway Platoon
The environment set $\mathcal{S}$ of a highway platoon consists of two lanes, going from South to North. We randomize initial locations of agents on two lanes. Agents aim to complete the trip quickly with action set $A_{i}$ for agent $i$. Meanwhile, there will be some background vehicles that have no driving strategies. The background vehicles are included in $\mathcal{S}$. Also, agents need to observe the position of other vehicles and make appropriate decisions to avoid accidents. The observation is included in $O_{i}$ for agent $i$.
Table II shows the set-up of agent $i$ at time $t$. In additional to the intersection set-up, the agent now can also choose to switch to the other lane without moving forward. Accordingly, we assigned a platoon reward of -2.
$s$ [${}_{i}[t]\in$ $\mathcal{S}$ Location and cumulative reward
$o$ [${}_{i}[t]\in$ $\mathcal{O}$ Observation of the current state space
$a$ [${}_{i}[t]\in$ $\mathcal{A}$ Go; Stop; Lane change: switch to the other lane
[…]
TABLE II: Platoon set-up
Figure 2: Highway Platoon
@@ III Methodology
Our study employs the OpenAI Chat API, specifically the GPT-4.0 model, to conduct our experiments. Notably, GPT-4.0 has demonstrated significant improvement over its predecessor, GPT-3.5, by at least 20%, as per OpenAI’s assertions, on an array of benchmarks such as the SuperGLUE language comprehension and the LAMBADA linguistic modeling tests. [33] As an advanced version of the transformer-based language model, GPT-4.0 holds considerable potential for traffic planning and autonomous driving.
III-A Prompt Design
@@
The chat model facilitates conversations through three distinct roles, including ’user’, ’assistant’, and ’system’. ’User’ represents the human participant interacting with the AI. The ’assistant’ role embodies the AI providing responses to the user’s prompts. Finally, the ’system’ role is primarily responsible for defining the assistant’s behavior—this can involve setting a specific language, character, or viewpoint for the assistant. The chat model prompt is structured as a series of messages, each identified by the role of the sender and the message content. By concatenating past interactions from both the user and assistant, the model can more effectively engage in multi-turn conversations.
In our research, we tailored our experiment without considering any past states. This led us to use only one ”system” message and one ”user” message to construct our prompt. The ”system” message outlines the general set-up, defining parameters like the environment space $\mathcal{S}$ (i.e., the road layout), action space $\mathcal{A}$ (i.e., all the permitted moves), and the reward function $\mathcal{R}$. Meanwhile, the ”user” message takes the perspective of a specific car. This message first identifies agent $i$ (i.e., its color) and its present observation $o_{i}$ (i.e., the locations of all the cars and its own cumulative reward). After that, the prompt asks the agent about the action $a_{i}$ it should take and the resulting location due to the action of choice.
Figure 3: Prompt and response structure
III-B Programming
We selected Python as our programming language and utilized the chat model to choose the move and calculate the location. Our programming was responsible for managing the progression of the game and computing the ground truth reward.
@@ III-B Programming
Our focus was on studying self-driven decision-making rather than cooperative decision-making. Therefore, we took precautions to ensure that when an agent made a decision, it did not have access to the concurrent decisions made by others. We did not ask the chat model to simulate the game in one prompted response, as it has access to the private observations of all agents and can implicitly use other agents’ private observations to make decisions. Instead, we simulated the traffic in a turn-based manner, where each turn represents a time step. Each inquiry to the chat model could only take the perspective of one agent. To further prevent access to concurrent decisions made by other agents, the game progression was updated at the end of each time step, after receiving responses from all agents.
With the above considerations in mind, we designed our program as shown in Figure 4. The program ran a loop through multiple time steps. At each time step, the program invoked the chat model from the perspective of each individual agent. After receiving responses from all agents, the code updated the reward and each agent’s current location. Subsequently, the game state was also updated. When a car completed its trip, we removed it from the game, meaning that it was no longer observed by other cars, prompted about its actions, and its location and reward stopped updating. This prevented any potential interference with the cars still in the game. The entire game’s progress from the beginning to the end was exported into a .txt file and a .cvs file.
Figure 4: Program structure (update picture later)
@@ IV Experiments
We modeled the environments of both the intersection and platoon scenarios using a grid consisting of multiple cells, with each cell assigned a pair of coordinates. In our experiments, we only consider one green car, one red car, and one white car. The red and green cars are participating agents, while the white car serves as a background vehicle. The goal of the participating agents is to complete the trip and maximize their cumulative reward by choosing various actions. The background vehicle’s goal is simply to complete the trip, with its policy dictating that it will choose ”Go” at each time step.
@@
The set-up is stored in a configuration file for ease of adjustment. This file contains all prompts, a list of cars, and their initial positions. Given that the general set-up is consistent for all agents and is unaffected by time, the system message portion of the prompt is stored as a static string. On the other hand, each car’s private observation, influenced by time and specific agent conditions, is stored as templates. Each time these templates are accessed, they are customized in accordance with present conditions.
As previously mentioned, each car could observe both the location and the cumulative reward of itself, but could only see the locations (not the cumulative rewards) of all other cars. To easily concatenate the observation about all other cars, we stored the user message as two separate templates: one for the observing itself, and one for all other cars. We assembled all these components of the prompt together in our programming.
@@
In both scenarios, we utilized the OpenAI API with the specified parameters: model = gpt-4, temperature = 0.1, max_token = 10. Increasing the temperature enhances the randomness of the response. [34] Our testing revealed that raising the temperature beyond 1.0 sometimes resulted in responses that did not conform to the desired format. Further increasing the temperature to 2.0 led to responses that were essentially unintelligible, as exemplified in Figure 5. The max_tokens parameter restricts the maximum length of the response. As per OpenAI’s tokenizer, the language model can consistently provide responses in the preferred format within 10 tokens. [35]
Figure 5: OpenAI confidently provides an unintelligible response (temperature = 2.0)
IV-A Scenario 1: Unsignalized Intersection
We modeled the environment of intersection using a 9 by 9 grid. Figure 6 shows an example of a prompt and a response. We adjust the prompt according to the number and placement of the background vehicles.
In order to inspect the behavior of LLMs, we decided to simulate the game multiple times. We simulated the Markov Game 50 times without BVs(short for background vehicles), 10 times with 1 to 4 BVs respectively. The starting positions of 4 BVs are (5,2),(5,3),(4,5),(5,5). Since we set background vehicles as non-stopping white cars, we have to allocate them this way to avoid crashing on background vehicles.
We want to explicitly investigate the behavior of Green car and Red car when there is no interference of background vehicles. We found out that Red car tends to stop before several grids of the intersection, so we define the stops at positions other than (4,5) or (5,4) as early stops. Figure 7 is the line plot of the frequency of early stops among each simulation with crashed simulations represented by blue dotted line.
From the plot, we know that Green car never stops earlier until it reaches the intersection and Red car stops when it is more than 1 grid away from the intersection. We can possibly infer that the LLM of Red car tends to make cautious decisions and behave courteously. Now, we can take a look at the regular stops meaning that they stop at (4,5) or (5,4). Figure 8 is the line plot of the frequency of regular stops among each simulation with crashed simulations represented by blue dotted line.
Figure 6: Intersection prompt and response example
Number of BVs Number of Simulations Number of Successes Success Rate
@@ IV-A Scenario 1: Unsignalized Intersection
1 10 10 1
2 10 10 1
3 10 10 1
4 10 10 1
TABLE III: Unsignalized Intersection Results
Figure 7: Frequency of Early Stops
Figure 8: Frequency of Regular Stops
From the plot, we know that both car tends to make multiple stop before the intersection regardless the other car’s position. We can possibly infer that the both cars wait long enough and one of them makes an action $a_{i}$=Go at timestamp i to finish the game. Now, we can investigate what happens in the crashed simulations which are 6-th and 37-th simulation. In 6-th simulation, Red car didn’t stop earlier and decided to Go at the same time as Green car so that crash happened. In 37-th simulation, Red car stopped earlier and Green car stopped regularly before the intersection. After Red car arrived at intersection, they both began to wait and again decided to Go at the same time.
Given the context of Markov Game, they have a 96% success rate. When they stop simultaneously, we believe that it mimics the social norm in human society. The random Go action mimics the driver who breaks the dilemma. However, human drivers could always stop if they notice the other driver chose to Go. That is the scenario that we can’t simulate by Markov Game.
IV-B Scenario 2: Highway Platoon
We modeled the environment of intersection using a 2 by 9 grid. Figure 9 shows an example of a prompt and a response. We adjust the prompt according to the number and placement of the background vehicles.
In this scenario, we did similar thing from previous scenario. We simulated the Markov Game 50 times without BVs, 10 times with 1 to 4 BVs respectively. The starting positions of 4 BVs are (1,2), (2,2), (1,4), (2,4). With our parameters set up, we find that in this scenario, the LLM performed excellently, and there were no car crashes in the simulations.
In this scenario, our focus is primarily on the frequency and timing of lane changes, which are crucial for platoon formation. Additionally, we perform an analysis of the scenario in the absence of interference from background vehicles. Figure 10 is the line plot of the frequency of lane changes in each simulation.
From the plot, we can see that there is at most one lane change per simulation, with the green car executing most of the lane changes. This is reasonable, as it suggests that agents are striving to minimize lane changes, which are a significant factor in car crashes on highways. Furthermore, the LLM of the green car may exhibit a more aggressive or consistent lane-changing behavior.
Figure 9: Platoon prompt and response example
Number of BVs Number of Simulations Number of Successes Success Rate

### v2

@@ Abstract
Social norm is defined as a shared standard of acceptable behavior in a society. The emergence of social norms fosters coordination among agents without any hard-coded rules, which is crucial for the large-scale deployment of autonomous vehicles (AVs) in an intelligent transportation system. This paper explores the application of large language models (LLMs) in understanding and modeling social norms in autonomous driving games. We introduce LLMs into autonomous driving games as intelligent agents who make decisions according to text prompts. These agents are referred to as LLM agents. Our framework involves LLM agents playing Markov games in a multi-agent system (MAS), allowing us to investigate the emergence of social norms among individual agents. We aim to identify social norms by designing prompts and utilizing LLMs on textual information related to the environment setup and the observations of LLM agents. Using the OpenAI Chat API powered by GPT-4.0, we conduct experiments to simulate interactions and evaluate the performance of LLM agents in two driving scenarios: unsignalized intersection and highway platoon. The results show that LLM agents can handle dynamically changing environments in Markov games, and social norms evolve among LLM agents in both scenarios. In the intersection game, LLM agents tend to adopt a conservative driving policy when facing a potential car crash. The advantage of LLM agents in games lies in their strong operability and analyzability, which facilitate experimental design.
@@ I-B Contributions of this paper
This work focuses on how LLMs can facilitate navigation in autonomous driving games. Our study utilizes the OpenAI Chat API [33], powered by GPT-4.0, to conduct experiments. Our contributions include:
@@
    We propose a framework where LLM agents play Markov games in a multi-agent system (MAS) in order to investigate whether social norms emerge among individual agents.
@@
    We design prompts to simulate interactions in autonomous driving games. The prompts consist of “system” messages outlining the general settings (e.g., the road layout) and “user” messages regarding the observations and decision making of LLM agents.
@@
    We apply the proposed framework to two traffic scenarios: the unsignalized intersection and highway platoon, to examine social norms formed by LLM agents.
The rest of the paper is organized as follows: In Section II, we first present preliminaries regarding Markov game, LLMs and social norm. Section III introduces our framework in which LLM agents play Markov games via prompt design. Section IV presents numerical experiments and results. Section V concludes.
Figure 1: LLM for autonomous driving games
@@ II-A Markov Game
A Markov game, also known as a stochastic game, is a mathematical framework that generalizes both Markov Decision Processes (MDPs) and game theory to model dynamic interactions among rational players in uncertain environments. In its general form, it is defined by a state set $\mathcal{S}$ and action sets $\mathcal{A}_{1},...,\mathcal{A}_{n}$ for each corresponding agent in the environment [34]. Each agent also has a corresponding reward function $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ for agent $i$. The goal of the Markov Game is to find an optimal policy for each agent. In this paper, we investigate the behavior and decision-making of LLMs when they play the Markov Game as agents. We use $\mathcal{S}$ to denote the overall environment, $\mathcal{A}_{i}$ to denote the action set of agent $i$, $\mathcal{O}_{i}$ to denote the observation set of agent $i$, $\mathcal{R}_{i}$ to denote the current reward of agent $i$. LLM chooses one of the actions from $\mathcal{A}_{i}$, with $\mathcal{O}_{i}$ and partial $\mathcal{S}$ as input. In this paper, the action choosing policy of an agent was solely decided by LLMs. Our primary focus is on two games as follows:
Scenario 1-Unsignalized intersection: The environment $\mathcal{S}$ is visualized in Figure 2(a), with Road 1 extending from the West to the East and Road 2 extending from the North to the South. The green and red cars represent the agents traveling on Road 1 and Road 2, respectively. Agents aim to cross the intersection, reach the end of the roads, and maximize their cumulative rewards. Table I summarizes the setup of agent $i$ at the beginning of time $t$. The location of agent $i$, $x_{i,t}$ corresponds to the cell it currently occupies. At each time step, each agent can choose ”Go” or ”Stop”, each with different action reward. The observation $o_{i,t}$ is a subset of current state space $\mathcal{S}$. Background vehicles (marked in white) with predetermined driving strategies are non-strategic players in the driving environment. The game terminates when all agents complete their trips or a crash occurs.
Scenario 2-Highway platoon: The environment $\mathcal{S}$ consists of two lanes, shows in Figure 2(b), going from North to South. Red and green agents each occupy one lane. Agents aim to reach the end of the lanes and maximize their cumulative rewards. In addition to the ”Go” and ”Stop” action, agents can now also choose to switch lanes. We defined the success of the game as forming a platoon at least once and completing the trips without any car crashes. A platoon is characterized by the presence of cars in the same lane. With the incentive of the reward and the overall goal, the agents will aim to complete their trips as quickly as possible, while avoiding accidents and attempting to maintain platoons. Table I summarize the setup of the platoon game.
((a)) Scenario 1
((b)) Scenario 2
Figure 2: Game scenarios
Scenario 1
$s_{i,t}\in\mathcal{S}$ Location $x_{i,t}$ and cumulative reward $\rho_{i,t}$
$o_{i,t}\in\mathcal{O}$ Observation of the current environment
$a_{i,t}\in\mathcal{A}$ Go: move forward by one cell; Stop: no move
$r_{i,t}\in\mathcal{R}$ Go: -2; Stop: -2; Crash: -5
Scenario 2
$s_{i,t}\in\mathcal{S}$ Location $x_{i,t}$ and cumulative reward $\rho_{i,t}$
$o_{i,t}\in\mathcal{O}$ Observation of the current state space
$a_{i,t}\in\mathcal{A}$ Go; Stop; Lane change: switch to the other lane
$r_{i,t}\in\mathcal{R}$ Go; Stop; Lane: -4; Crash: -5; Platoon: +2
TABLE I: The setup of game scenarios
II-B Large Language Models
@@
A large language model (LLM) is a type of artificial intelligence algorithm designed to understand, generate, and interact with human language at a large scale [33]. These models are trained on vast amounts of text data, learning patterns, vocabulary, grammar, and even nuances of language such as context, tone, and implications. They use a specific architecture known as Transformer [35], which allows them to efficiently process sequences of words and predict the next word in a sentence, understand the meaning of a text, or generate new text that follows a given prompt [36]. LLMs are highly versatile AI tools with applications spanning multiple industries and fields [37]. They excel in tasks such as content creation, customer support, education, language translation, software development, legal documentation, healthcare research, optimization, sentiment analysis, and accessibility technologies. In this paper, we examine the application of LLMs in autonomous driving games to see if how they make decisions and form social norms.
II-C Social Norms
Social norms represent the collective expectations regarding appropriate conduct within groups [38]. They encompass both the unwritten agreements that dictate how society members should act, and the formalized regulations and laws that are established. Social norms in car driving [32] are the unwritten rules that guide driver behavior on the road. Our primary focus is on examining whether expected social norms can evolve among LLM agents. The expected social norm in Scenario 1 is “yielding to others” at the intersection. The expected social norm in Scenario 2 is the formation of a platoon among agents.
@@ III Methodology
In this section, we present the framework of leveraging LLMs to understand social norms in Markov games.
In Figure 1, we utilize prompt-chaining for LLM interactions in the Markov game. At each time step, the program invoked the chat model from the perspective of each individual agent. The games are submitted to LLMs as prompts in which the respective game, including the choice options, is described. Once LLMs have made their choices (i.e., the actions made by strategic and non-strategic agents), which we track as a completion of the given text, we update the prompts and then submit the new prompt to LLMs for the next state in the game. The textual information includes the reward and each agent’s current location. When a car completes its trip, we remove it from the game, meaning that it is no longer observed by other cars, or prompted about its actions. Its location and reward stop updating as well. This prevents any potential interference with the cars still in the game. The output includes the decision making, rewards and states of agents in the Markov game, which is used to analyze social norms.
@@
Prompt design: The chat model facilitates conversations through three distinct roles, including “user”, “assistant”, and “system” (See Figure 3). “User” represents the human participant interacting with the AI. The “assistant” role embodies the AI providing responses to the user’s prompts. Finally, the “system” role is primarily responsible for defining the assistant’s behavior—this can involve setting a specific language, character, or viewpoint for the assistant. The chat model prompt is structured as a series of messages, each identified by the role of the sender and the message content. By concatenating past interactions from both the user and assistant, the model can more effectively engage in multi-turn conversations. The detailed prompts are in Section IV. In our research, we assume that the agents will not be capable of making predictions based on any historical context regarding previous states. This leads us to use only one “system” message and one “user” message to construct our prompt. The “system” message outlines the general settings, defining parameters like the environment space $\mathcal{S}$ (i.e., the road layout), action space $\mathcal{A}$ (i.e., all the permitted moves), and the reward function $\mathcal{R}$. Meanwhile, the “user” message takes the perspective of a specific car. This message first identifies agent $i$ (i.e., its color) and its present observation $o_{i}$ (i.e., the locations of all the cars and its own cumulative reward). After that, the prompt asks the agent about the action $a_{i}$ it should take and the resulting location due to the action of choice.
Figure 3: Prompt design
We employ the OpenAI Chat API, specifically the GPT-4.0 model, to conduct our experiments. Notably, GPT-4.0 has demonstrated significant improvement over its predecessor, GPT-3.5, by at least 20%, as per OpenAI’s assertions, on an array of benchmarks such as the SuperGLUE language comprehension and the LAMBADA linguistic modeling tests. As an advanced version of the transformer-based language model [33], GPT-4.0 holds considerable potential for traffic planning and autonomous driving.
@@
Note that our focus is not cooperative game. Therefore, we ensure that when an agent made a decision, it does not have access to the concurrent decisions made by others. We do not ask the chat model to simulate the entire game in one prompted response, as it has access to the private observations of all agents and can implicitly use other agents’ private observations to make decisions. Instead, we simulate the traffic in a turn-based manner, where each turn represents a time step. Each inquiry to the chat model could only take the perspective of one agent. To further prevent access to concurrent decisions made by other agents, the game state is updated at the end of each time step, after receiving responses from all agents.
@@ IV Experiments
In this section, we conduct experiments on two game scenarios to understand how social norms can be formed when LLMs play Markov games.
@@
Experiment setup: The environments of both the intersection and platoon scenarios adopt a grid representation consisting of multiple cells, with each cell assigned a pair of coordinates to denote the location of vehicles. The red and green cars are strategic agents, while the white car serves as a background vehicle. All background vehicles are not involved in the formation of social norms. The setup is stored in a configuration containing all prompts, a list of cars, and their initial positions. Given that the general settings is consistent for all agents and is unaffected by time, the system message portion of the prompt is stored as a static string. On the other hand, each car’s private observation, influenced by time and specific agent conditions, is stored as templates. Each time these templates are accessed, they are customized in accordance with present conditions. Each car could observe both the location and the cumulative reward of itself, but could only see the locations of all other cars. To easily concatenate the observation about all other cars, we store the user message as two separate templates: one for the observing itself, and one for all other cars.
@@
In both scenarios, we utilize the OpenAI API with the specified parameters: model = gpt-4, temperature = 0.1, max token = 10. Increasing the temperature enhances the randomness of the response [39]. Our testing reveal that raising the temperature beyond 1.0 sometimes results in unstable behaviors. Further increasing the temperature to 2.0 leads to responses that are essentially unintelligible, as exemplified in Figure 4. The max token parameter restricts the maximum length of the response. As per OpenAI’s tokenizer, the language model can consistently provide responses in the preferred format within 10 tokens [40].
Figure 4: Responses (temperature = 2.0)
IV-A Scenario 1-Unsignalized intersection
Figure 5: Prompts and responses - Scenario 1
# of BVs # of tests # of Yields Rate
@@ IV-A Scenario 1-Unsignalized intersection
1 50 50 1
2 50 50 1
3 50 50 1
4 50 50 1
TABLE II: Performance - unsignalized intersection
Figure 5 shows an example of our prompts for scenario where the environment of an unsignalized intersection has a 9 by 9 grid. In order to inspect the behavior of LLMs, we simulate the Markov Game 50 times. Table II summarizes the performance of LLMs when varying numbers of background vehicles. The results demonstrate the formation of social norms at the unsignalized intersection (i.e., yielding to others). It is noteworthy that crashes may occur when no background vehicles are present. This can be attributed to the fact that the presence of background vehicles restricts the action space for ego cars, thereby reducing the likelihood of vehicles encounters at the intersection and facing a crash.
We now look into the case when there are no background vehicles. The results show that the red car tends to early stop before several grids of the intersection. The early stop behavior measures the waiting time of vehicles. It is defined as the stop action of vehicles when they haven’t arrived at the intersection (5,5), i.e., vehicles stop at positions other than (4,5) or (5,4), 1 grid away from the intersection. Figure 6 visualizes the line plot of the number of early stops in each simulation. The x-axis represents the simulation. The y-axis represents the number of early stops in each simulation. The red and green lines denote the early stop behaviors of red and green vehicles, respectively. The crash cases are marked by blue dotted lines. The average number of early stops of the red car is 1.42, and the average of early stops of green car is 0.0. The results show that a conservative driving policy is adopted by the red car when playing the Markov game at the unsignalized intersection.
We then investigate the crash cases, which occurr in the 6th and 37th simulations. In the 6th simulation, the red car does not stop earlier and decides to proceed at the same time as the green car, resulting in a crash. In the 37th simulation, the red car stops earlier, while the green car stops 1 grid away from the intersection. When the red car arrives at the intersection, both cars decide to proceed simultaneously, leading to a crash.
Figure 6: Early stop behaviors
To eliminate the early stop behaviors and shorten the waiting time of vehicles, we increase the reward of Go action from -2 to 0. We re-simulate the Markov game for 50 times. The results show that (1) The average number of early stops for the red car decreases from 1.42 to 0.48, suggesting that the reward design can influence driving behavior. (2) The success rate of adherence to social norms decreases from 0.96 to 0.94 due to an increase in the number of car crashes. It indicates a trade-off between waiting time and road safety in the intersection game. The decrease in waiting time would increase the possibility of crashes.
IV-B Scenario 2-Highway Platoon
Figure 7 shows an example of prompts for scenario where the environment of a highway platoon has a 2 by 9 grid.
Figure 7: Prompts and responses - Scenario 2
# of BVs # of tests # of Platoons Rate